OpenStudy (anonymous):

(n 3) 2 · (n5)4 find the product

Parth (parthkohli):

\( \color{Black}{\Rightarrow\large (n^3)^2 \cdot (n^5)^4 }\) This?

OpenStudy (anonymous):

yes! can you help me please

Parth (parthkohli):

We'll have to use these two laws: \( \color{Black}{\Rightarrow \Large (a^x)^y = a^{xy} }\) and \( \color{Black}{\Rightarrow \Large a^x \cdot a^y = a^{xy} }\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow \large n^{3\cdot 2} \cdot n^{5 \cdot 4} }\)

Parth (parthkohli):

@moneec24 reply?

OpenStudy (anonymous):

OK